OpenStudy (anonymous):

WILL GIVE MEDAL. c: Solve for x: -3|x + 7| = -12

OpenStudy (tkhunny):

Have you considered dividing by -3?

OpenStudy (anonymous):

no, i literally have no clue how to do this, i'm trying this class out before taking it next year

OpenStudy (tkhunny):

Can you solve this? 3x = 15

OpenStudy (anonymous):

divide 3x by 3 and 15 by 3 which gives you 5, right?

OpenStudy (anonymous):

3x/3 will give me x

OpenStudy (anonymous):

x=5

OpenStudy (tkhunny):

Very good. Why is the present problem any different? Divide by -3, EXACTLY like you just divided by 3 in my little sample problem.

OpenStudy (anonymous):

i divide -3 by -12?

OpenStudy (anonymous):

the |x+7| thing confuses me

OpenStudy (anonymous):

i divided -3 by -12 and i got 0.25, is that right?

OpenStudy (tkhunny):

Why does it matter? 3x = 12 ==> x = 4 3Frog = 12 ==> Frog = 4 3yz = 12 ==> yz = 4 3(x+2) = 12 ==> (x+2) = 4 3|x+7| = 12 ==> |x+7| = 4 3Pittsburgh = 12 ==> Pittsburgh = 4 Works every time. You are dividing by -3. You are not dividing by -12.

OpenStudy (anonymous):

ohhhh, i think i got it

OpenStudy (anonymous):

can you help with another please?

OpenStudy (tkhunny):

Post new question on a new thread. gtg

OpenStudy (anonymous):

okay ty